# FP-Inconsistent

### v1

@@ 4.2 Anti-bot services
We integrated two popular commercial anti-bot services on our honey site: DataDome [29] and BotD [30]. Both DataDome and BotD provide real-time decisions on requests received on a website. DataDome advertises real-time decision for a request in under 2 milliseconds with an overall accuracy of 99% and a false positive rate of 0.01%. BotD is based on the popular open-source fingerprinting library FingerprintJS [31]. BotD claims using “the most advanced device fingerprinting technology”, and reports a detection accuracy of 99.5%.
@@
We integrate JavaScript libraries of both these services on our honey site111 As required by DataDome, for each request, we also make an API call from our server to get their decision. These libraries collect information about the visiting user on the website and relay this information to their own servers. The servers respond with the final decision of whether a real human or a bot originated the request.
@@
These services are black-boxed and do not provide information on the browser attributes they use as features to decide if a request originates from a bot. To determine this information, we crawl our honey site using OpenWPM [32]. OpenWPM is an open source tool which can be used to track the behavior and interaction of different web elements, including scripts, on a webpage.
Table V in Appendix A highlights the different browser APIs used by DataDome and BotD. We find that both these services make use of a number of popular fingerprinting APIs such as HTMLCanvasElement.getContext (allows drawing on the browser canvas), window.navigator.userAgent (returns the browser user agent), window.navigator.plugins (returns the plugins supported by the browser) and more. There are a few APIs which are only used by DataDome, such as window.navigator.serviceWorker (provides information about whether the browser supports service workers), window.navigator.maxTouchPoints (maximum number of simultaneous touch points supported by the device), and window.navigator.hardwareConcurrency (number of logical processors available on the device).
On the other hand, there are only two APIs which are accessed by BotD and not by DataDome: window.navigator.productSub (returns the build number of the browser), and window.navigator.appVersion (returns version information for the browser). Overall, we find that DataDome collects more attributes from each request as compared to BotD. In later sections, we see that DataDome has higher accuracy in detecting bots over BotD. The additional attributes collected by DataDome potentially help better detection.
@@
Figure 2: Screenshot from a seller on SEOClerks making claims about sending organic traffic to drive engagement on websites. The claims likely suggest that the seller employs bots that employ evasive fingerprints to took real users.
4.3 Bot services
@@
We made purchases from multiple online bot services to send traffic to different versions of our honey site. We make most of or purchases from the SEOClerks [12], an underground marketplace for web traffic where bot services advertise their traffic as being real, organic and adsense safe to boost website engagement. Their claims of being able to send real and organic traffic indicate that they are likely using evasive bots that alter their fingerprints to look like real users. Figure 2 captures a screenshot from a bot service on SEOClerks making such claims about their traffic. In addition to purchases from SEOClerks, we also purchase traffic from Babylon Traffic [11] and Spark Traffic [13]. While making these purchases we share URLs with different version strings with different bot services to identify the bot services of each request on our honey site.
4.4 Data Collected
Figure 3: Overview of our data collection pipeline
@@ 4.4 Data Collected
To characterize the differences in the browser fingerprints of evasive bots, we extract information from different browser APIs and properties upon loading our honey site in the browser. We send send this information to our server in an http request. We use FingerprintJS [31], a widely deployed browser fingerprinting library to capture this information. FingerprintJS captures over 30 different browser attributes including the list of fonts installed on the browser, the number of CPU cores on the device running the browser, the amount of memory on the device running the browser, the languages supported by the browser. While we predominantly focus on the attributes captured by FingerprintJS in our measurement analysis, we also capture features computed by CreepJS [33] along with keyboard and mouse interactions of the user.
5 Measurement analysis
@@
TABLE I: Overview of different bot services sending requests to our honey site including their evasion rate against DataDome and BotD.
@@
In this section, we analyze the requests obtained on our honey site. Over a period of 3 months, we received a total of 507,080 requests from 20 different bot services. We first report the detection rate of the anti-bot services and then compare browser attributes of bots that evade detection against those that were detected. This analysis helps understand the attributes used by bots to evade detection and ways to overcome them.
@@
Table I shows the statistics of the requests obtained from each bot service along with the evasion rate against the two anti-bot services on our honey site (DataDome and BotD). Among the 507,080 requests we received, 55.44% of requests were detected by DataDome and 47.07% of requests were detected by BotD. These results show that a significant proportion of bots are able to look like real users and evade anti-bot services.
Takeaway 1: Our measurement shows concrete evidence of evasive bots that are not detected by commercial anti-bot services. This shows that any characterization of differences in fingerprints between bots and real users based on the decisions of bot detection is unreliable since there exist bots that are able to present themselves as real users.
Figure 4: Pie chart showing the proportion of requests originating from IP addresses with ASNs that are not present in block lists, ASNs that are present in block lists but evading and detected by DataDome.
Figure 5: Pie chart showing the proportion of requests originating from IP addresses with ASNs that are not present in block lists, ASNs that are present in block lists but evading and detected by BotD.
5.1 IP addresses for evasion
@@
Among the requests received on our honey site, we observed requests from IP addresses with Autonomous System Numbers (ASNs) mapping to cloud services such as Amazon Web Services (AWS), DigitalOcean etc. Since such ASNs are likely flagged as those used by bots [34, 35], we check the ASNs of the requests we received against public ASN block lists [36, 37]. Figure 4 and Figure 5 show that 82.54% of requests originated from flagged ASNs. Among these, 52.93% of requests evade BotD and 43.17% of requests evade BotD. These results show that evasive bots are able to evade detection even when they send requests from flagged ASNs.
@@
We suspect that anti-bot services may not rely on ASN block lists to reduce false positives since real users and bots can share the same ASNs but can send requests from different IP addresses. Accordingly, we ran similar analysis with blocked IP addresses using MaxMind’s minFraud API [38]. Similar to findings in prior research [19], we find that IP block lists offer limited coverage (15.86%). However, among the IP addresses that were covered, requests from 48.1% were able to evade DataDome and 68.85% were able to evade BotD. We present detailed analysis in the Appendix D.
@@
In conclusion, we see that a significant number of bots that sent requests from blocked IP addresses and ASNs were able to evade both DataDome and BotD. This indicates that evasive bots change request attributes and don’t merely send requests from IP requests not captured by block lists to evade detection.
@@
Takeaway 2: Evasive bots change request attributes and don’t merely rely on sending requests from IP addresses that are not captured by block lists to evade detection.
5.2 Browser attributes for evasion
@@
Since evasive bots don’t merely rely on IP addresses, we systematically analyze the values of browser attributes to identify those used by evasive bots for evasion. Concretely, we train models to distinguish between the requests that were detected by and evaded DataDome and BotD respectively. We then use techniques from explainability of machine learning to identify browser attribute values that help with evasion. We then explore the values of these attributes on requests from bot services that were most successful with evasion to verify that they enable evasion.
5.2.1 Identifying browser attributes
@@
We train two random forest classifiers using XGBoost [39] to distinguish between the requests that were detected and evaded DataDome and BotD respectively. Each classifier takes feature vectors representing browser attributes in each request (discussed in Section 4.4) as input and provides a binary decision on whether that request would detected by the respective anti-bot service.
@@
We performed a 90-10 split on the requests we received to train the classifiers. The classifier for BotD attained an accuracy 97.8% on the training set and 97.71% on the test set while the classifier for DataDome attained an accuracy of 82.09% on the training set and 81.66% on the test set. These high values for accuracy indicate that the browser attributes of requests that evade the two anti-bot services are considerably different from those of requests detected by them.
@@
TABLE II: Top 5 most important browser attributes that help evade DataDome and BotD
@@
We use SHapley Additive exPlanations or SHAP [40] to analyze the decisions made by these classifiers to identify browser attributes that result in evasion. Table Table II lists the top 5 features that help evade DataDome and BotD respectively.
5.3 Browser attributes among evasive bots
@@
We now inspect the attribute values of requests from bot services that have high evasion rates to see if they exploit the previously identified attributes for evasion. Concretely, we compare attribute values across bot services that have high evasion rate against those that have low evasion rates against the anti-bot services.
@@
Figure 6: Bar plot showing the probability of PDF plugins that have the highest probability of evasion against BotD. This plot shows a weakness of BotD where the presence of any plugin helps evade BotD
5.3.1 Bots evading BotD
@@
We inspected requests from the top 3 bot services with highest evasion rates against BotD (S15, S18 and S19 in Table I) and the top 3 bot services with lowest evasion rates against BotD (S6, S16 and S17 in I). We record 7,132 requests from the top 3 bot services evading BotD and report 100% evasion among them. We record 39,620 requests from the top 3 bot services that are detected by BotD and report an evasion rate of 5.11% among them.
@@
We did not observe significant differences between the values of vendor flavors, vendor and touch support attributes among requests from these bot services. 99.91% of requests from services evading BotD supported the Chrome PDF Viewer plugin, while 100% of requests detected by BotD did not support any plugins. Motivated by these stark differences, we further investigate the impact of plugins on evading BotD. Concretely, from all requests received on our honey site, we compute the probability of evading BotD when supporting any one of 5 commonly used PDF plugins. Figure 6 shows that the presence of any PDF plugin nearly guarantees evasion against BotD.
Takeaway 3: Plugins are a potential blind spot for BotD. Evasive bots are able to evade BotD by showing support for at least one standard PDF plugin.
5.3.2 Bots evading DataDome
@@
We similarly inspect requests from the top 3 bot services with highest evasion rates against DataDome. We record 52,746 requests from the top 3 bot services evading DataDome (S8, S9 and S17 in Table I) and report 79.15% evasion among them. We record 51,110 requests from the top 3 bot services that are detected by DataDome (S7, S11 and S16 in Table I) and report an evasion rate of 4.12%.
@@
Figure 7: Cumulative probability distribution function (CDF) plots of the number of CPU cores recorded on requests from bot services that had the highest evasion rate over DataDome against those that had the the lowest evasion rate over DataDome.
@@
100% of requests from the top 3 bot services having the highest evasion rate against DataDome did not support any plugins. However, not using plugins does not ensure evasion against DataDome since 56.45% of requests from the 3 bot services with the lowest detection rate against DataDome did not support any plugins either. Analyzing screen frame and forced colors attributes across these bot services that evade and are detected by DataDome revealed certain values that always result in detection. However, we did not observe values for these attributes that help with evasion.
@@
Figure 7 compares cumulative probability distribution functions (CDFs) of the values for the number of cores (captured by hardwareConcurrency) on requests from bot services with high evasion rates over DataDome against the values on requests from bot services with low evasion rates over DataDome. These results indicate that low values for hardwareConcurrency help evade DataDome. Concretely, we see that 84.7% of requests from bot services with high evasion rate against DataDome used devices having fewer than 8 cores. In contrast, 38.16% of requests from bot services detected by DataDome used devices with fewer than 8 cores. We can further assess the impact of hardwareConcurrency by disregarding requests that contain values for screen frame and forced colors attributes that always lead to evasion. With this refinement, we see that 84.7% of requests from bot services with high evasion rate against DataDome employ devices with fewer than 8 cores while only 19.05% of requests from bot services with low evasion rate against DataDome employ devices with fewer than 8 cores.
@@
Thus, we observe that evasive bots evading DataDome ensure certain values being presents for combinations of attributes. This is different from evasive bots evading BotD that ensured certain values for one set of features (plugins). We investigate more combinations of attributes that help evade DataDome in Appendix E.
Takeaway 4: While hardwareConcurrency is a potential blind spot for DataDome, merely ensuring low values for hardwareConcurrency is not sufficient for evasion. Evasive bots are able to evade DataDome by ensuring low values for hardwareConcurrency and making sure that other attributes don’t lead to detection.
5.3.3 Bots evading DataDome and BotD
@@
Among the requests received on our honey site, the requests from two different bot services have over 80% evasion rate against both DataDome and BotD (S14 and S20 in Table I). We receive a total of 5,302 requests from these two bot services which have 84.7% evasion rate against DataDome and 90.59% evasion against BotD.
@@
We observe that 83.77% of these requests have fewer than 8 CPU cores indicating that they exploit hardware concurrency to evade DataDome. Interestingly, 93.02% of these requests do not have any plugins, indicating that they do not exploit BotD’s blind spot towards plugins for evasion. These requests exploit touchSupport, a different blind spot of BotD for evasion. Concretely, 78.36% of requests from the bot services evading both DataDome and BotD support touch events, while only 3.95% of requests from the top 3 bot services having the lowest evasion rate against BotD support touch events. In contrast, only 0.07% of requests from the top 3 bot services that only evaded BotD (Section 5.3.1) showed support for touch events and 8.61% of requests from the top 3 bot services that only evaded DataDome (Section 5.3.2) showed support for touch events.
Takeaway 5: touchSupport is another blind spot for BotD. Some evasive bots indicate support for touch events to evade BotD instead of supporting plugins.
6 Inconsistency analysis

### v2

@@ 4.2. Anti-bot services
We integrated two popular commercial anti-bot services on our honey site: DataDome (DataDome, [n. d.]) and BotD (Fingerprint, [n. d.]). Both DataDome and BotD provide real-time decisions on requests received on a website. DataDome advertises real-time decisions for a request in under 2 milliseconds with an overall accuracy of 99% and a false positive rate of 0.01%. Prior research on bot detection has explored DataDome (Amin Azad et al., 2020; Yasuhara et al., 2024). BotD is a bot detection service from the developers of the popular open-source fingerprinting library FingerprintJS (Fingerprint, [n. d.]) that is widely used in industry and academia (Kalantari et al., 2024; Li et al., 2021; Amin Azad et al., 2020; Vastel et al., 2020). BotD claims to use “the most advanced device fingerprinting technology”, and reports a detection accuracy of 99.5%.
@@
We integrate JavaScript libraries of both these services on our honey site111 As required by DataDome, for each request, we also make an API call from our server to get their decision. These libraries collect browser fingerprints of the browser visiting the honey site and relay them to their own servers. The servers then respond with the decision of whether a real human or a bot originated the request.
@@
These services are black-boxed and do not provide information on fingerprint attributes they use as features to decide if a request originates from a bot. To determine this information, we crawl our honey site using OpenWPM (OpenWPM, [n. d.]). OpenWPM is an open-source tool to track the behavior of different web elements, including scripts, on a webpage.
Table 5 in Appendix B highlights the different browser APIs accessed by DataDome and BotD. Both services access a number of fingerprinting APIs such as navigator.plugins, HTMLCanvasElement.getContext, navigator.userAgent, and more. We find that DataDome collects more attributes from each request than BotD. In Section 5, we see that DataDome has higher bot detection accuracy than BotD, which could potentially stem from these additional attributes.
@@
Figure 2. Screenshot from a bot service on SEOClerks making claims about sending organic traffic to drive engagement on websites. The claims likely suggest that the bot service employs evasive bots to took real users.
4.3. Bot services
@@
We made purchases from multiple online bot services to send traffic to different versions of our honey site. We make our purchases from the SEOClerks (seoclerks, [n. d.]), an underground marketplace for web traffic where bot services advertise their traffic as being real, organic, and Adsense safe to boost website engagement. Their claims of being able to send real and organic traffic indicate that they are likely using evasive bots that alter their fingerprints to look like real users. Figure 2 captures a screenshot from a bot service on SEOClerks making such claims about their traffic. We share URLs with different version strings with different bot services to identify the bot services of each request on our honey site.
4.4. Data Collection
Figure 3. Overview of our data collection pipeline.
@@ 4.4. Data Collection
To characterize the differences in the fingerprint attributes of evasive bots, we extract information from different browser APIs and properties upon loading our honey site in the browser. We send this information to our server in an http request. We use FingerprintJS (Fingerprint, [n. d.]), a widely deployed browser fingerprinting library to capture this information. FingerprintJS captures over 30 different fingerprint attributes including the list of fonts installed on the browser, the number of CPU cores on the device running the browser, the amount of memory on the device, and the languages supported by the browser. While we focus on the attributes captured by FingerprintJS in this paper, both our measurement analysis (Section 5) and our methodology to discover inconsistencies (Section 7) are compatible with other fingerprint attributes too.
5. Analysis
@@
Table 1. Overview of different bot services sending traffic to our honey site and their evasion rates against DataDome and BotD.
@@
Over a period of 3 months, from September 2023 to November 2023, we received 507,080 requests from 20 different bot services. We first report the detection rate of the anti-bot services and then compare fingerprint attributes of bots that evade detection against those that were detected. This analysis helps understand the attributes used by bots for evasion and ways to overcome them.
@@
Table 1 shows the statistics of the traffic obtained from each bot service along with the evasion rate against the two anti-bot services on our honey site (DataDome and BotD). Among the 507,080 requests we received, 55.44% of requests were detected by DataDome, and 47.07% of requests were detected by BotD. These results show that a significant proportion of bots are able to evade anti-bot services.
Takeaway 1: Our measurement shows that evasive bots are not reliably detected by commercial anti-bot services.
5.1. IP addresses for evasion
@@
We observed requests on our honey site that contained IP addresses with Autonomous System Numbers (ASNs) mapping to cloud services such as Amazon Web Services (AWS). Since such ASNs are likely flagged as those used by bots (Chen et al., 2013; Hu et al., 2009), we check the ASNs of the requests we received against public ASN block lists (brianhama, [n. d.]; growtoups, [n. d.]). We report that 82.54% of requests originated from flagged ASNs. Among these, 52.93% of requests evade BotD and 43.17% of requests evade BotD. These results show that evasive bots are able to evade detection even when they send requests from flagged ASNs.
@@
We suspect that anti-bot services may not rely on ASN block lists since real users and bots can share the same ASNs but can send requests from different IP addresses. Accordingly, we ran similar analysis with blocked IP addresses using MaxMind’s minFraud API (MaxMind, 2024b). Consistent with findings in prior research (Li et al., 2021), we find that IP block lists offer limited coverage (15.86%). More interestingly, among the IP addresses that were covered, requests from 48.1% were able to evade DataDome and 68.85% were able to evade BotD.
@@
In conclusion, we see that a significant number of bots that sent requests from blocked IP addresses and ASNs were able to evade both DataDome and BotD. This indicates that evasive bots don’t merely send requests from IP addresses not captured by block lists to evade detection.
@@
Takeaway 2: Evasive bots do not merely rely on sending requests from IP addresses that are not captured by block lists to evade detection.
5.2. Fingerprint attributes for evasion
@@
Since evasive bots don’t merely rely on IP addresses, we systematically analyze the browser fingerprint attributes in their requests to identify those used for evasion. Concretely, we train models to distinguish between the requests that were detected by and evaded DataDome and BotD respectively. We then use techniques from the explainability of machine learning to identify fingerprint attribute values that help with evasion. We then explore the values of these attributes on requests from bot services that were most successful with evasion to verify that they enable evasion.
5.2.1. Identifying fingerprint attributes
@@
We train two random forest classifiers using XGBoost (XGBoost, [n. d.]) to distinguish between the requests that were detected and evaded DataDome and BotD respectively. Each classifier takes as input fingerprint attributes from each request (discussed in Section 4.4) and provides a binary decision on whether that request would detected by the respective anti-bot service.
@@
We performed a 90-10 split on the requests to train the classifiers. The classifier for BotD attained an accuracy 97.8% on the training set and 97.71% on the test set while the classifier for DataDome attained an accuracy of 82.09% on the training set and 81.66% on the test set. These high accuracy values indicate that the fingerprint attributes of requests that evade the two anti-bot services are considerably different from those of requests detected by them.
@@
Table 2. Top 5 most important fingerprint attributes that help evade DataDome and BotD.
@@
We use SHapley Additive exPlanations or SHAP (SHapley Additive exPlanations, [n. d.]) to analyze these classifiers to identify fingerprint attributes that result in evasion. Table Table 2 lists the top 5 attributes that help evade DataDome and BotD respectively.
5.3. Fingerprint attributes among evasive bots
@@
We now inspect the attribute values of requests from bot services with high evasion rates to see if they exploit the attributes identified in Table 2 for evasion. Concretely, we compare attribute values across bot services that have high evasion rates against those that have low evasion rates.
@@
Figure 4. Bar plot showing the probability of PDF plugins that have the highest probability of evasion against BotD. This plot shows that the presence of any plugin helps evade BotD.
5.3.1. Bots evading BotD
@@
We inspected requests from the top 3 bot services with the highest evasion rates against BotD (S15, S18, and S19 in Table 1) and the top 3 bot services with the lowest evasion rates against BotD (S6, S16, and S17 in 1). We record 7,132 requests from the top 3 bot services evading BotD and report 100% evasion among them. We record 39,620 requests from the top 3 bot services that are detected by BotD and report an evasion rate of 5.11% among them.
@@
We did not observe significant differences between the values of Vendor Flavors, Vendor, and Touch Support attributes among requests from these bot services. 99.91% of requests from services evading BotD supported the Chrome PDF Viewer plugin, while 100% of requests detected by BotD did not support any plugins. Motivated by these stark differences, we further investigate the impact of plugins on evading BotD. Concretely, from all requests received on our honey site, we compute the probability of evading BotD when supporting any one of 5 commonly used PDF plugins. Figure 4 shows that the presence of any PDF plugin nearly guarantees evasion against BotD.
5.3.2. Bots evading DataDome
@@
We similarly inspect requests from the top 3 bot services with the highest and lowest evasion rates against DataDome. We record 52,746 requests from the top 3 bot services evading DataDome (S8, S9, and S17 in Table 1) having 79.15% evasion among them. We record 51,110 requests from the top 3 bot services detected by DataDome (S7, S11, and S16 in Table 1) with an evasion rate of 4.12%.
@@
Figure 5. Cumulative probability distribution function (CDF) plots of the number of CPU cores recorded on requests from bot services that had the highest evasion rate over DataDome against those that had the the lowest evasion rate over DataDome.
@@
100% of requests from the top 3 bot services having the highest evasion rate against DataDome did not support any plugins. However, 56.45% of requests from the 3 bot services with the lowest evasion rate against DataDome did not support any plugins either. Analyzing the Screen Frame and Forced Colors attributes revealed certain values that always result in detection. However, we did not observe values for these attributes that help with evasion.
@@
Figure 5 compares cumulative probability distribution functions (CDFs) of the number of CPU cores (captured by hardwareConcurrency) on requests from bot services with high evasion rates over DataDome against the values on requests from bot services with low evasion rates over DataDome. These results indicate that low values for hardwareConcurrency help evade DataDome. Concretely, 84.7% of requests from bot services with a high evasion rate against DataDome had fewer than 8 cores. In contrast, only 38.16% of requests from bot services detected by DataDome had fewer than 8 cores. To further assess the impact of hardwareConcurrency, we disregard requests that contain values for Screen Frame and Forced Colors that always lead to evasion. Now, 84.7% of requests from bot services with a high evasion rate against DataDome have fewer than 8 cores while only 19.05% of requests from bot services with a low evasion rate against DataDome have fewer than 8 cores.
@@
Evasive bots evading DataDome ensure certain values for combinations of attributes. This is different from evasive bots evading BotD that ensured certain values for one set of attributes (plugins). We investigate more combinations of attributes that help evade DataDome in Appendix C.
5.3.3. Bots evading DataDome and BotD
@@
Requests from two different bot services have over 80% evasion rate against both DataDome and BotD (S14 and S20 in Table 1). We received 5,302 requests from these services which have an 84.7% evasion rate against DataDome and 90.59% evasion against BotD.
@@
We observe that 83.77% of these requests have fewer than 8 CPU cores indicating that they exploit hardware concurrency to evade DataDome. Interestingly, 93.02% of these requests do not have any plugins, indicating that they do not exploit plugins to evade BotD. They exploit touchSupport, a different blind spot of BotD for evasion. Concretely, 78.36% of requests from the bot services evading both DataDome and BotD support touch events, while only 3.95% of requests from the top 3 bot services having the lowest evasion rate against BotD support touch events. In contrast, only 0.07% of requests from the top 3 bot services that only evaded BotD (Section 5.3.1) showed support for touch events and 8.61% of requests from the top 3 bot services that only evaded DataDome (Section 5.3.2) showed support for touch events.
Takeaway 3: Evasive bots exploit either touchSupport and plugins to evade BotD. They exploit hardwareConcurrency to evade DataDome.
6. Inconsistency analysis